OpenStudy (lcnettech):

PLZ HELP will medal.........What is the surface area of this rectangular prism?

OpenStudy (lcnettech):

418cm2 520cm2 836cm2 209cm2

OpenStudy (butterflydreamer):

do you know how to find surface area?

OpenStudy (lcnettech):

nope

OpenStudy (butterflydreamer):

what about area?

OpenStudy (lcnettech):

yes

OpenStudy (lcnettech):

base times hieght

OpenStudy (butterflydreamer):

yep! Excellent :). So when they ask you for surface area it means : The TOTAL area of the surface of a 3-Dimensional object.

OpenStudy (lcnettech):

so 8 times 5?

OpenStudy (lcnettech):

or 13 times 5

OpenStudy (butterflydreamer):

Firstly, how many sides does the rectangular prism have?

OpenStudy (lcnettech):

6

OpenStudy (butterflydreamer):

Yes! So that means you have to find the area of the 6 sides and then add them all together to get the TOTAL area

OpenStudy (butterflydreamer):

do you understand?

OpenStudy (lcnettech):

i got 520 when i did 5*8*13.do i times that by 2?

OpenStudy (butterflydreamer):

Not quite. If you do 5*8*13 that will give you volume and not surface area. You have to find the area of each side and add them together.

OpenStudy (lcnettech):

13*13.....5*5....and 8*8 ?is that what i do?

OpenStudy (butterflydreamer):

So let's begin with the smallest sides! How do you find the area of the smallest rectangle?

OpenStudy (lcnettech):

you do 5*5?

OpenStudy (butterflydreamer):

area is base times height, remember? So the smallest rectangle has a base of 8 and a height of 5 right?

OpenStudy (butterflydreamer):

What would the area of that rectangle be?

OpenStudy (lcnettech):

40

OpenStudy (butterflydreamer):

yep! So 40 is just for that one small rectangle. But remember, there is another small rectangle which is the same size. So what would you do?

OpenStudy (lcnettech):

do 40 *2?

OpenStudy (lcnettech):

which is 80

OpenStudy (butterflydreamer):

Exactly! So we know that the total surface area of the smallest rectangles = 80 Now, we do the same for the other rectangles.

OpenStudy (lcnettech):

13*5=65*2=130

OpenStudy (butterflydreamer):

yep :) You're understanding this really well. Okay and now you have one final set of rectangles to go.

OpenStudy (lcnettech):

is it 8 x 13?

OpenStudy (butterflydreamer):

yeah but there are two rectangles that size! (Top and bottom)

OpenStudy (lcnettech):

8*13=104......104*2=208

OpenStudy (butterflydreamer):

Yes! So now the TOTAL surface area would be?

OpenStudy (lcnettech):

418

OpenStudy (butterflydreamer):

YAY :D! Great work .

OpenStudy (lcnettech):

thank you so much. :D

OpenStudy (butterflydreamer):

No worries :) hehe

OpenStudy (lcnettech):

would this formula be the same for a triangular prism?

OpenStudy (butterflydreamer):

yep :) Just remember that for surface area of any prism, you just add all the areas together. So in a triangular prism, you find the area of the two triangles and then the other three rectangle sides. Add it all up, and you should get your answer!

OpenStudy (lcnettech):

ok thanks:)

OpenStudy (butterflydreamer):

you're welcome !!

OpenStudy (lcnettech):

let me know if you need help i any thing.sorry it took so long im just bad with surface area.

OpenStudy (butterflydreamer):

Sure will and the same goes for you! haha don't worry, i was horrible at it also. Just practice :)

OpenStudy (anonymous):

surface area=2(l*b+b*h+h*l)

OpenStudy (butterflydreamer):

LOL that was so much simpler... xD But yes, @lcnettech you can also just use the equation provided by @surjithayer . Surface area = 2(l*b + b*h + h*l)